OpenStudy (anonymous):

Evaluate 2x-4y for x = 2 and y = 4.

OpenStudy (anonymous):

substitute.. 2(2) - 4(4) = 4 - 16 = -12

OpenStudy (anonymous):

A. -64 B. 1/8 C. 1/2

OpenStudy (anonymous):

hmm w8 a sec

OpenStudy (anonymous):

Um, Ok?

OpenStudy (anonymous):

sorry its 1/2

OpenStudy (anonymous):

Ok thanks (:

OpenStudy (anonymous):

2*4/2^(-4)=8/16=1/2